OpenStudy (anonymous):

If four angles of a convex pentagon measure 120°, 93°, 108°, and 101°, find the measure of the fifth angle. Show equations and work that lead to your answer.

OpenStudy (anonymous):

sum of angles in triangle = (3-2)*180 = 180 sum of angles in quadrilateral = (4-2)*180 = 360 and so on with the pattern: sum of angles in n-sides shape = (n-2)*180 so for the pentagon, calculate sum of its angles, then you can find the fifth angle

OpenStudy (anonymous):

118 degree

OpenStudy (anonymous):

okay thanks I just got it!:)